OpenStudy (anonymous):

I've got a differential equation: e^x*y*y'=e^(-y)+e^(-2x-y) I found an implicit solution: y^2(e^y)=-4e^(-x)-(4/3)e^(-3x)+c Could you guys check this? If this is right, is there a way to get an explicit solution?

OpenStudy (anonymous):

e^(x)(y)(y')

OpenStudy (anonymous):

I would say: \[e^xyy'=e^{-y}+e^{-2x}e^{-y} \implies e^xy \frac{dy}{dx}=e^{-y}(1+e^{-2x})\] \[e^yydy=e^{-x}(1+e^{-2x})dx\] Integrating: \[\int\limits y e^y dy=\int\limits (e^{-x}+e^{-3x})dx \implies ye^y-\int\limits e^ydy=-e^{-x}-\frac{1}{3}e^{-3x}+C \implies ye^y-e^y=\]\[-e^{-x}-\frac{1}{3}e^{-3x}+C\]

OpenStudy (anonymous):

What did you do?

OpenStudy (anonymous):

I got (e^y)(y^2)/4 on the left side, and multiplied through by 4.

OpenStudy (anonymous):

Ok, I did the wrong integration by parts.

OpenStudy (anonymous):

I did mine in my head so mine may be wrong D: